How many grams of KBr are contained in 200 mL of a 0.310 M KBr solution?

Please show ALL of your work!
Answer #1

Answer

first we will find the number of moles of KBr by the molarity formula

moles of solute(KBr)= molarity*volume of the solution in L

moles of KBr= 0.310*0.2=0.062 moles of KBr

now finding the mass of KBr by the formula

mass= moles*molar mass

mass of KBr= moles of KBr*molar mass of KBr

mass of KBr= 0.062*119.002 = 7.378 grams of KBr